Basic Salary of an employee is Rs. 18,000 and Allowances are Rs. 9000. According to the company’s policy casual leaves are 12 days per year and earned leaves are 24 days per year and normal working days are 22 per month.

Find the total cost of leaves as percent of gross salary.
1
Expert's answer
2019-05-20T03:09:27-0400

Total number of leaves is 12+24 =36 days. Total number of working days in year is 12*22 = 264. If every day of leave is payed as usual working day, then total cost of leaves as percent of gross salary is 36/264*100% = 3*12/(22*12)*100% =

3/22*100% = 13.636363%

Answer: 13.636363%.